OpenStudy (anonymous):

Which is an equivalent expression to negative thirty-three plus negative twenty-four? negative thirty-three minus positive twenty-four positive thirty-three minus positive twenty-four negative twenty-four plus positive thirty-three positive twenty-four minus negative thirty-three

OpenStudy (anonymous):

@texaschic101

OpenStudy (cggurumanjunath):

negative thirty-three =? @iceicebaby

OpenStudy (anonymous):

-33

OpenStudy (cggurumanjunath):

good

OpenStudy (anonymous):

what??

OpenStudy (cggurumanjunath):

negative twenty-four =? @iceicebaby

OpenStudy (anonymous):

-24

OpenStudy (cggurumanjunath):

very good

OpenStudy (cggurumanjunath):

-33+(-24) = ? @iceicebaby

OpenStudy (anonymous):

idk

OpenStudy (cggurumanjunath):

-33-24=? @iceicebaby

OpenStudy (cggurumanjunath):

-(33+24) =? @iceicebaby

OpenStudy (cggurumanjunath):

negative thirty-three minus positive twenty-four

OpenStudy (cggurumanjunath):

negative thirty-three =? @iceicebaby

OpenStudy (anonymous):

chill hold up

OpenStudy (cggurumanjunath):

coolllllllllllllllllllllllllllllllllllllllllllllllllllllllllllllllllllllllllllllllllllllll

OpenStudy (cggurumanjunath):

@iceicebaby